# PyTomography: A Python Library for Medical Image Reconstruction Luke Polson1,2, Roberto Fedrigo1,2, Chenguang Li1,2, Maziar Sabouri1,2, Obed Dzikunu2,3, Shadab Ahamed1,2, Nicolas Karakatsanis4, Arman Rahmim1,2,3,5,∗, Carlos Uribe2,5,6,∗

### v3

@@ Abstract
Background: There is an absence of open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction algorithms available in recent literature, such as those that employ artificial intelligence.
Purpose: To create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms.
Methods: PyTomography was developed using Python and inherits the GPU-accelerated functionality of PyTorch and parallelproj for fast computations. Its flexible and modular design decouples the system matrix from reconstruction algorithms, simplifying the process of integrating new imaging modalities using various python tools.
Results: Example use cases demonstrate the software capabilities in parallel hole SPECT and listmode PET imaging. Use of the OSEM, BSREM, KEM, and DIPRecon reconstruction algorithms are explored. System matrices incorporating breathing-motion correction are explored in SPECT imaging. PyTomography is validated against alternative vendor/open source software for OSEM based reconstruction.
Conclusions: We have developed and publicly shared PyTomography, a highly optimized and user-friendly software for medical image reconstruction, with a class hierarchy that fosters the development of novel imaging applications.
I. Introduction
@@
Medical imaging forms a cornerstone in modern healthcare by providing visual and quantitative information about internal body structures and functions. It enables early detection1, accurate diagnosis, and precise treatment planning for a wide range of medical conditions2, 3, 4, contributing to improved patient outcomes and enhanced medical decision-making. Reconstruction algorithms are routinely used to generate 3D images that can be used for both research and clinical decision making 5.
@@
While the development and validation of reconstruction techniques remains an active field of research6, 7, 8, 9, it can often be difficult to share and disseminate one’s findings. Additionally, while manufacturers of tomographic imaging equipment such as Single Photon Emission Computed Tomography (SPECT) and Positron Emission Tomography (PET) cameras provide their own internal reconstruction software, users cannot access all implementation details; this limits reproducibility in scientific studies. These issues could be solved by moving to an open source paradigm for medical image reconstruction. Here, the imaging community could openly collaborate in the development of reconstruction techniques. Adoption of a singular open-source framework, as opposed to private frameworks developed independently by various research groups, would strengthen the ability to develop, share, and compare findings. This could ultimately help accelerate the development and translation of new diagnostic imaging capabilities into clinical practice.
@@
A few tools have already been developed to try and address some of these issues. STIR 10 and CASToR 11 are two C/C++ based open source and collaborative image reconstruction frameworks that support a variety of reconstruction algorithms for SPECT and PET. NiftyPET 12 is a python library which employs GPU-accelerated code written in CUDA C for PET reconstruction. QSPECT13 and TIRIUS support use of the ordered subset expectation maximum (OSEM) algorithm for SPECT and PET reconstruction respectively. Other libraries, such as TomoPy 14 and the ASTRA toolbox 15, support a variety of GPU-accelerated image tomographic reconstruction algorithms, although they lack extensive SPECT/PET system modeling. Web links to all these tools, as well as a list of other available tools for image reconstruction, can be found on the NMMItools webpage of the Society of Nuclear Medicine and Molecular Imaging (SNMMI). The unique purpose of this research is to build upon recent efforts in GPU-accelerated reconstruction16, 17, 18, 19 and AI-based reconstruction7, 8, 9 to establish an open source medical image reconstruction programming framework written in a popular programming language that facilitates collaboration. Valuable past efforts have focused on CUDA programming to accelerate reconstruction times in SPECT16 and PET 16, 17 reconstruction through the use of multi-threading. However, potential contributors to an open-source framework may be unfamiliar with CUDA; any tool that prioritizes community contributions should employ familiar backend code. In Python, the library NumPy is widely known, but it lacks direct GPU acceleration support. PyTorch, however, emerges as a suitable alternative since it maintains a similar interface to NumPy while also facilitating GPU acceleration and direct interface with neural networks. Some potential contributors may wish to provide contributions using GPU-accelerated functionality offered by other libraries, such as CuPy; an optimal framework should thus also permit the ability to incorporate alternative implementations for particular tasks.
@@
To this end, the python library PyTomography has been developed with the following priorities: (i) implementation of standard and traditional imaging modalities and reconstruction algorithms, (ii) implementation and dissemination of recent research developments, such as the deep image prior8, and (iii) encouragement of community involvement via extensive documentation and user-friendly tutorials. While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT/PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
@@
Overall, the aim of this study is to introduce the community to the PyTomography package, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development. To demonstrate the versatility of the software, example applications will be explored in (i) SPECT reconstruction using PyTorch based rotate+sum projectors, and (ii) listmode (LM) time of flight (TOF) PET reconstruction using Joseph’s projectors from the python library parallelproj19. In what follows, we elaborate on our methods and results for PyTomography, which is made publicly available to download; relevant links can be found in section VI..
II. Materials and Methods
II.A. Mathematical Description
@@
A brief overview of the mathematical paradigm of medical imaging is presented in order to motivate the software architecture. The mathematical conventions used in this paper are as follows. Vectors are represented by lower case letters, while linear operators are represented by upper case letters. The product of two vectors $vw$ and division of two vectors $v/w$ are inferred to be point-wise operations. Components of a vector are given by $v_{i}$, and components of linear operators are given by $A_{ij}$.
II.A.1. Projectors
@@
In the paradigm of tomographic medical imaging, there are two main vector spaces: the “projection” space $\mathcal{V}$ consisting of measured projection data $g$ and the “object” space $\mathcal{U}$ consisting of 3D objects $f$. Since measured and reconstructed data is digitized, $\mathcal{U}$ and $\mathcal{V}$ are assumed here to be finite dimensional vector spaces. In SPECT imaging, for example, $\mathcal{V}$ consists of a set of 2D pixelated images acquired by counting photons emitted by radiopharmaceuticals, and $\mathcal{U}$ represents the voxelized spatial distribution of radioactivity concentration in the object being imaged.
@@
Certain imaging systems can be characterized by a system matrix $H$, which corresponds to a linear operator mapping from object space to projection space. Its components $H_{ij}$ represent a contribution of voxel $j$ to detector element $i$. In SPECT/PET, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ represents scatters and randoms. In CT, analytical modeling is achieved via $\bar{g}=be^{-Hf}$ where $b$ denotes the rate of emissions from the transmission source to a particular measurement. The operation of $H$ on an object vector is known as forward projection (FP), while the operation of $H^{T}$ on a projection vector is known as back projection (BP).
Efficient implementation of FP and BP depends on the particular imaging modality; this typically involves decomposing $H$ into a sequence of less computationally burdensome operations. In general, FP can be decomposed into the following sequence: (i) linear endomorphisms mapping from $\mathcal{U}\to\mathcal{U}$, (ii) operators mappings from $\mathcal{U}\to\mathcal{V}$, and (iii) linear endomorphisms mapping from $\mathcal{V}\to\mathcal{V}$ given by $\{B_{k}\}$. BP is given by the transpose of this operation sequence. In PyTomography, operators which map to the same vector space are defined as transforms, while operators that map between different vector spaces are defined as projectors. Example implementations available in PyTomography are listed below:
@@
  • Parallel Collimator SPECT: Modeling of primary detections in this imaging modality can be approximated using the “rotate+sum” technique16, whereby (i) consists of rotations and attenuation/collimator detector response modeling, (ii) consists of summation along an axis, and (iii) consists of adjustments for a camera with a finite field of view. System modeling is approximated using the following system matrix:
@@
    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ (transform) implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ (transform) implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ (transform) implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ (transform) implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ (projector) is summation along the $+x$ axis, and $\mathcal{T}_{\text{fov}}^{(n)}$ (transform) applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
  • LM PET: Modeling in this modality can be approximated when (i) consists of object-space resolution modeling, (ii) consists of ray tracing between all detected lines of response (LOR), and (iii) consists of normalization/attenuation correction for each LOR. In this case, the system matrix can be written as
@@
    where $\mathcal{T}_{rr}$ (transform) is used for object-space resolution modeling, $\mathcal{P}_{\Sigma}$ (projector) is used to compute line integrals, $\mathcal{T}_{\eta}$ (transform) adjusts for detector normalization, and $\mathcal{T}_{\omega}$ (transform) adjusts for attenuation effects. Due to cancellation in certain reconstruction algorithms, $\mathcal{T}_{\eta}$ and $\mathcal{T}_{\omega}$ need not always be included (though they are required for computing the sensitivity factor).
@@
where $\left\{A_{n}\right\}$ are object-to-object space transforms, $\left\{H_{n}\right\}$ are a sequence of system matrices, $\left\{B_{n}\right\}$ are a sequence of projection-to-projection space transforms, $v_{n}$ is a basis vector of length $N$ with a value of $1$ in component $n$ and $\otimes$ is an outer product. This can be utilized in the following examples:
  • Periodic Motion Correction: $A_{n}$ represent spatial transformation operators corresponding to time gate $n$, $H_{n}$ represent system matrices for each motion transform (attenuation modeling may depend on object deformation), and $B_{n}=1$. This maps to an extended projection space $(N,...)$ where $N$ is the number of different time gates (e.g. during a breathing cycle).
@@
II.A.2. Likelihoods
A likelihood function $L(g|\bar{g})$ provides the likelihood of observing data $g$ given expectation $\bar{g}$. Since $\bar{g}$ is approximated via analytical models depending on $f$ and $H$, it follows that $L(g|\bar{g})=L(g|f,H,...)$, where the ellipsis “...” contains additional parameters such as a scatter estimate $s$; the ellipsis will assumed implicitly for the remainder of this paper. Many iterative reconstruction algorithms seek to maximize the likelihood function via gradient ascent using the gradient $\nabla_{f}L(g|f,H)$. Likelihood functions for different modalities are shown in Table 1.
Table 1: Likelihoods and corresponding gradients for various medical imaging modalities; $s$ represents a scatter estimate used in SPECT/PET while $b$ represents a scaling factor used in CT.
@@ II.A.2. Likelihoods
II.A.3. Reconstruction Algorithms
An image reconstruction algorithm $A$ uses measured data $g$ (i.e. projections) to estimate a corresponding object $f$ that would produce projections $g$ given a system matrix $H$. Statistical algorithms find the image estimate $f$ which maximizes a likelihood function $L(g|H,f)$ characterizing the imaging system. This can be expressed as
@@ II.A.3. Reconstruction Algorithms
$f^{n+1}=f^{n}+C^{n}(f^{n},H_{n})\left[\nabla_{f}L(g^{n}|f^{n},H_{n})-\beta_{n}\nabla_{f}V(f^{n})\right]$ (5)
where $n$ is the subiteration index, $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H_{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta_{n}=c_{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c_{n}$). These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once. Various PGA algorithms available in PyTomography are shown in Table 2; they are characterized by their preconditioner factor. Additional reconstruction techniques can be developed using these algorithms. The kernel expectation maximum (KEM), for example, corresponds to (i) reconstructing a coefficient image $\alpha$ using $\alpha=A_{\text{OSEM}}(g,HK)$ where $K$ is a kernel transform and (ii) obtaining the image estimate via $f=K\alpha$.
Table 2: Sample algorithms available in PyTomography: shown are ordered subset expectation maximum (OSEM), ordered subset maximum a posteriori one step late (OSMAPOSL), block sequential regualized expectation maximum (BSREM), and kernelized expectation maximum (KEM).
Algorithm Preconditioner factor $C^{n}(f^{n},H)$
OSEM20 $f^{n}/\left(H_{n}^{T}1\right)$ and $V(f^{n})=0$
OSMAPOSL21 $f^{n}/\left(H_{n}^{T}1+\nabla_{f}V(f^{n})\right)$
BSREM22 $f^{n}\alpha(n)/\left(\omega_{n}H^{T}1\right)$
Certain algorithm names in the literature correspond to specific realizations of the generic algorithms in Table 2 with particular likelihoods in Table 1. The OS-SART algorithm, for example, corresponds to the use of OSEM with the weighted MSE likelihood (with weights $W=1/(H1)$).
@@
where $w_{rs}$ provides a weighting factor between the locations voxel $r$ and voxel $s$ and $\phi(f_{r},f_{s})$ characterizes a weighting factor between the observed image values at voxel $r$ and voxel $s$. When $w_{rs}$ is non-zero only for neighbouring voxels, the prior is known as a nearest neighbour prior; various nearest neighbour priors implemented in PyTomography are shown in Table 3.
@@
Relative difference penalty (RDP) 24 $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it is also possible to allow it to depend on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT6. In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map.
PyTomography also offers support for the Deep Image Prior Reconstruction (DIPRecon) algorithm. Details regarding the implementation of this algorithm can be found in Algorithm 2 of Gong et. al8. The purpose of this reconstruction technique is to use anatomical images (CT/MR) to enhance functional images (SPECT/PET) through a sequence of neural network overfitting and mixing with OSEM updates. This reconstruction algorithm requires the use of an anatomical image $z$ and a neural network architecture $\mathcal{N}(\cdot;\theta)$ that can be trained to predict a reconstructed functional image $f$ given $z$ as input. This paper explores the use of DIPRecon on an ultra-high resolution PET/MR brain phantom.
II.B. Software Architecture
The key points of the last section are that (i) medical imaging consists of many different modalities that each require unique implementations for system modeling and (ii) tomographic reconstruction algorithms are modality agnostic. PyTomography has been carefully designed to permit arbitrary implementations of system modeling that can interface with the available reconstruction algorithms. The components of PyTomography, designed to facilitate the architecture described, are listed below.

### v4

@@ Abstract
There is a need for open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction algorithms available in recent literature, such as those that employ artificial intelligence. The purpose of this research was to create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms. PyTomography was developed using Python and inherits the GPU-accelerated functionality of PyTorch and parallelproj for fast computations. Its flexible and modular design decouples system matrices, likelihoods, and reconstruction algorithms, simplifying the process of integrating new imaging modalities using various python tools. Example use cases demonstrate the software capabilities in parallel hole SPECT and listmode PET imaging. Overall, we have developed and publicly share PyTomography, a highly optimized and user-friendly software for medical image reconstruction, with a class hierarchy that fosters the development of novel imaging applications.
keywords:
Medical Imaging , Image Reconstruction , SPECT , PET
journal: SoftwareX
Metadata
Nr. Code metadata description Please fill in this column
C1 Current code version v3.0.0
C2 Permanent link to code/repository used for this code version https://github.com/qurit/PyTomography
C3 Permanent link to Reproducible Capsule None. Reproducible examples provided as part of the documentation.
C4 Legal Code License MIT License
C5 Code versioning system used git
C6 Software code languages, tools, and services used python
C7 Compilation requirements, operating environments & dependencies python >=>=> = 3.11
C8 If available Link to developer documentation/manual https://pytomography.readthedocs.io/en/latest/
C9 Support email for questions https://pytomography.discourse.group/
Table 1: Code metadata (mandatory)
1 Motivation and significance
@@
Medical imaging forms a cornerstone in modern healthcare by providing visual and quantitative information about internal body structures and functions. It enables early detection [1], accurate diagnosis, and precise treatment planning for a wide range of medical conditions [2, 3, 4], contributing to improved patient outcomes and enhanced medical decision-making. Reconstruction algorithms are routinely used to generate 3D images that can be used for both research and clinical decision making [5].
@@
While the development and validation of reconstruction techniques remains an active field of research [6, 7, 8, 9], it can often be difficult to share and disseminate one’s findings. Additionally, while manufacturers of tomographic imaging equipment such as Single Photon Emission Computed Tomography (SPECT) and Positron Emission Tomography (PET) cameras provide their own internal reconstruction software, users cannot access all implementation details; this limits reproducibility in scientific studies. These issues could be solved by moving to an open source paradigm for medical image reconstruction. Here, the imaging community could openly collaborate in the development of reconstruction techniques. Adoption of a singular open-source framework, as opposed to private frameworks developed independently by various research groups, would strengthen the ability to develop, share, and compare findings. This could ultimately help accelerate the development and translation of new diagnostic imaging capabilities into clinical practice.
@@
A few tools have already been developed to try and address some of these issues. STIR [10] and CASToR [11] are two C/C++ based open source and collaborative image reconstruction frameworks that support a variety of reconstruction algorithms for SPECT and PET. NiftyPET [12] is a python library which employs GPU-accelerated code written in CUDA C for PET reconstruction. QSPECT [13] and TIRIUS support use of the ordered subset expectation maximum (OSEM) algorithm for SPECT and PET reconstruction respectively. Other libraries, such as TomoPy [14] and the ASTRA toolbox [15], support a variety of GPU-accelerated image tomographic reconstruction algorithms, although they lack extensive SPECT/PET system modeling. Web links to all these tools, as well as a list of other available tools for image reconstruction, can be found on the NMMItools webpage of the Society of Nuclear Medicine and Molecular Imaging (SNMMI).
The unique purpose of our present research and effort is to build upon recent efforts in GPU-accelerated reconstruction [16, 17, 18, 19] and AI-based reconstruction [7, 8, 9] to establish an open source medical image reconstruction programming framework written in a popular programming language that facilitates collaboration. Valuable past efforts have focused on CUDA programming in languages such as C/C++ to accelerate reconstruction times in SPECT [16] and PET [16, 17] reconstruction through the use of multi-threading. While some contributors to an open-source framework may prefer to use C/C++ programming, others may be unfamiliar with these languages; any tool that prioritizes community contributions should permit both options. Some of the advantages of a python framework are that (i) it is a widely known programming language and thus encourages community involvement and (ii) it permits the embedding of C/C++ code for particular applications. Python can directly utilize CUDA acceleration via the PyTorch library, for example, allowing users unfamiliar with C/C++ code to achieve GPU acceleration.
@@
To this end, we developed the python library PyTomography with the priorities to: (i) implement standard and traditional imaging modalities and reconstruction algorithms, (ii) implement and disseminate of recent research developments, such as the deep image prior [8], and (iii) encourage community involvement via extensive documentation and user-friendly tutorials. Based on the motivations previously described, the backend of library has been developed using PyTorch because it maintains a similar functional structure to the widely used NumPy library. Other libraries are used for more particular tasks; the python library paralleproj [19], which was developed with CUDA C backend code, is used for line integral computation in PET projectors. In general, code written in any language can be integrated in PyTomography provided a python interface is available. While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT and PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
@@
Overall, the aim of this study is to introduce the community to the PyTomography package, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development. To demonstrate the versatility of the software, example applications are explored in (i) SPECT reconstruction using PyTorch based rotate+sum projectors, and (ii) listmode (LM) time-of-flight (TOF) PET reconstruction. In what follows, we elaborate on our methods and results for PyTomography, which is made publicly available to download.
2 Software description
2.1 Software scope and functionality
@@
A brief overview of the mathematical paradigm of medical imaging is presented in order to motivate the software architecture. The mathematical conventions used in this paper are as follows. Vectors are represented by lower case letters, while linear operators are represented by upper case letters. The product of two vectors $vw$ and division of two vectors $v/w$ are inferred to be point-wise operations. Components of a vector are given by $v_{i}$, and components of linear operators are given by $A_{ij}$. When $A1$ or $A^{T}1$ is written, “1” is to be interpreted as a vector (in the appropriate vector space) with all elements equal to 1.
2.1.1 Projectors
@@
In the paradigm of tomographic medical imaging, there are two main vector spaces: the “projection” space $\mathcal{V}$ consisting of measured data $g$ and the “object” space $\mathcal{U}$ consisting of 3D objects $f$. Since measured and reconstructed data is digitized, $\mathcal{U}$ and $\mathcal{V}$ are assumed here to be finite dimensional vector spaces. In SPECT imaging, for example, $\mathcal{V}$ consists of a set of 2D pixelated images acquired by counting photons emitted by radiopharmaceuticals, and $\mathcal{U}$ represents the voxelized spatial distribution of radioactivity concentration in the object being imaged.
@@
Linear imaging systems can be characterized by a system matrix $H:\mathcal{U}\to\mathcal{V}$. Its components $H_{ij}$ represent the contribution of element $j$ in object space to element $i$ in projection space. In emission tomography, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ is an additive term containing estimates for scattered events (SPECT) or scatter and random coincidences (PET). In CT, analytical modeling is achieved via $\bar{g}=be^{-Hf}$ where $b$ denotes the rate of emissions from the transmission source to a particular measurement. The operation of $H$ on an object vector is known as forward projection (FP), while the operation of $H^{T}$ on a projection vector is known as back projection (BP). Example implementations available in PyTomography are listed below:
  • 1.
@@
    Parallel Collimator SPECT: Modeling of primary detections in this imaging modality can be approximated using the “rotate+sum” technique [16]; in this case, the system matrix can be written as
@@
    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ is summation along the $+x$ axis, and $\mathcal{T}_{\text{fov}}^{(n)}$ applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
  • 2.
    PET: Modeling of primary events (no randoms or scatters) in PET can be approximated as
@@
    where $\mathcal{T}_{rr}$ implements object-space resolution modeling (typically via a uniform Gaussian kernel), $\mathcal{P}_{\Sigma}$ implements line integral computation (listmode/sinogram), $\mathcal{T}_{\eta}$ adjusts for detector normalization, and $\mathcal{T}_{\omega}$ adjusts for attenuation effects.
@@
where $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ are additional operations, $\left\{H_{n}\right\}$ are a sequence of system matrices, $v_{n}$ is a basis vector of length $N$ with a value of $1$ in component $n$, and $\otimes$ is an outer product. This can be utilized in the following examples:
  • 1.
    Periodic Motion Correction: $A_{n}$ represent spatial deformation operators corresponding to time gate $n$, $H_{n}$ represent system matrices for each motion transform (attenuation modeling may depend on object deformation), and $B_{n}=1$. This maps to an extended projection space $(N,...)$ where $N$ is the number of different time gates (e.g. during a breathing cycle).
  • 2.
@@
2.1.2 Likelihoods
A statistical likelihood function $\tilde{L}(g|\bar{g})$ is proportional to the probability of observing an actual realization of data $g$ given expectation $\bar{g}$. Since $\bar{g}$ is approximated using $f$ and $H$, it follows that $\tilde{L}(g|\bar{g})=\tilde{L}(g|f,H,...)$, where the ellipsis “...” contains additional parameters such as a scatter estimate $s$; the ellipsis will be assumed implicit for the remainder of this paper. Many iterative reconstruction algorithms seek to maximize the likelihood function via gradient ascent by exploiting the gradient $\nabla_{f}\tilde{L}(g|f,H)$. Since (i) computation of $\nabla_{f}\log\tilde{L}(g|f,H)$ is more tenable than $\nabla_{f}\tilde{L}(g|f,H)$ and (ii) maximization of $\log\tilde{L}(g|f,H)$ is equivalent to maximization of $\tilde{L}(g|f,H)$, the statistical “log-likelihood” is nearly always considered instead. In PyTomography, the “likelihood” $L(g|f,H)$ refers to an objective function to be maximized; as such, the “likelihood” $L(g|f,H)$ is often used interchangeably with the statistical “log-likelihood” $\log\tilde{L}(g|f,H)$ for a particular modality. Likelihood functions for different modalities are shown in Table 2.
Table 2: Likelihoods and corresponding gradients for various medical imaging modalities; $s$ represents a scatter estimate used in SPECT/PET while $b$ represents a scaling factor used in CT. The CT likelihood corresponds to an ideal scenario where the number of counts in each sinogram bin follows a Poisson distribution; this is not necessarily the case in real systems with complex electronic noise.
@@ 2.1.2 Likelihoods
2.1.3 Reconstruction Algorithms
An image reconstruction algorithm $A$ uses measured data $g$ (i.e. projections) to estimate a corresponding object $f$ that would produce this $g$ given a system matrix $H$. Statistical algorithms find the image estimate $f$ which maximizes the likelihood function $L(g|H,f)$ characterizing the imaging system. This can be expressed as
@@ 2.1.3 Reconstruction Algorithms
$f^{n+1}=f^{n}+C^{n}(f^{n},H^{n})\left[\nabla_{f}L(g^{n}|f^{n},H^{n})-\beta^{n}\nabla_{f}V(f^{n})\right]$ (5)
where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H^{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta^{n}=c^{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{n}$). These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once. Various PGA algorithms available in PyTomography are shown in Table 3; they are characterized by their preconditioner factor. Certain algorithm names in the literature correspond to specific realizations of the generic algorithms in Table 3 with particular likelihoods in Table 2. The OS-SART algorithm, for example, corresponds to the use of OSEM with the weighted MSE likelihood (with weights $W=1/(H1)$).
Table 3: Sample algorithms available in PyTomography: shown are ordered subset expectation maximum (OSEM), ordered subset maximum a posteriori one step late (OSMAPOSL), block sequential regualized expectation maximum (BSREM), and kernelized expectation maximum (KEM). $n$ is used as a superscript (not an exponent) for all variables. For BSREM, $\alpha^{n}$ is a relaxation parameter.
Algorithm Preconditioner factor $C^{n}(f^{n},H^{n})$
OSEM [20] $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
OSMAPOSL [21] $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
BSREM [22] $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
@@
where $w_{rs}$ provides a weighting factor between the locations voxel $r$ and voxel $s$ and $\phi(f_{r},f_{s})$ characterizes a weighting factor between the observed image values at voxel $r$ and voxel $s$. When $w_{rs}$ is non-zero only for neighbouring voxels, the prior is known as a nearest neighbour prior; various nearest neighbour priors implemented in PyTomography are shown in Table 4. While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it is also possible to allow it to depend on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT [6]. In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map.
@@
Relative difference penalty (RDP) [23] $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
PyTomography also offers support for the Deep Image Prior Reconstruction (DIPRecon) algorithm. Details regarding the implementation of this algorithm can be found in Algorithm 2 of Gong et. al [8]. The purpose of this reconstruction technique is to use anatomical images (CT/MR) to enhance functional images (SPECT/PET) through a sequence of neural network overfitting and mixing with OSEM updates. This reconstruction algorithm requires the use of an anatomical image $z$ and a neural network architecture $\mathcal{N}(\cdot;\theta)$ that can be trained to predict a reconstructed functional image $f$ given $z$ as input. This paper explores the use of DIPRecon on an ultra-high resolution PET/MR brain phantom.
2.2 Software architecture
PyTomography has been carefully designed to permit arbitrary implementations of system matrices (projectors) which can be used in various likelihood functions that themselves interface with reconstruction algorithms. The components of PyTomography were designed to facilitate this architecture and are listed below. A visual representation of the software library along with sample code representing a SPECT reconstruction use case is shown in Figure 1.